OpenStudy (anonymous):

Use the distributive property to simplify the expression below: -4x-16/-4

OpenStudy (dape):

\[ \frac{-4x-16}{-4} = \frac{-4x}{-4}+\frac{-16}{-4} = x+4 \]

OpenStudy (anonymous):

thank you

OpenStudy (dape):

You're welcome.